# CodeSense: a Real-World Benchmark and Dataset for Code Semantic Reasoning

### v1

@@ 1 Introduction
Recent years have witnessed the emergence of numerous benchmarks for evaluating coding-related tasks. However, the majority of these efforts have focused on code generation using synthetic or narrowly scoped data—for example, HumanEval+ [1], LiveCodeBenchmark [2], Bigcodebench [3], and CodeBenchGen [4]—often extracted from isolated competitive programming problems. Consequently, they fail to capture the complexity and structure of real-world software development. Other benchmarks that incorporate real-world code, such as SWE-Bench [5], SWE-PolyBench [6], and KGym [7], tend to evaluate only task-specific performance (e.g., patch generation for GitHub issues), making it difficult to assess whether models exhibit generalizable semantic understanding. Finally, reasoning-focused benchmarks, such as CruxEval [8], primarily target function-level input/output prediction over short and synthetic code fragments involving random string operations. Such settings neglect the fine-grained semantic reasoning about internal program behavior and properties, data dependencies, and control structures required to solve a variety of SE tasks for complex real-world software systems.
To this end, we propose CodeSense, a benchmark for fine-grained code semantic reasoning, constructed from real-world GitHub projects in Python, C, and Java (see Table 1). CodeSense  introduces a spectrum of reasoning tasks at statement, code-block, and function levels, targeting essential semantic properties frequently needed across SE activities. For example, predicting loop iteration counts is critical for input/output prediction, performance analysis, and detecting infinite loops (e.g., denial-of-service vulnerabilities). Branch condition prediction and reasoning about pointers in C code are important for test input generation and memory safety assurance. As illustrated in Figures 1(a) and 1(b), fine-grained semantic reasoning about arithmetic operations, control flow, and API semantics is foundational for a variety of SE applications. Prior work [9, 10] has shown that incorporating semantic signals during training improves model performance on code generation, branch prediction, code clone detection, program repair and vulnerability detection tasks, motivating our design of CodeSense to comprehensively evaluate models’ capabilities for semantic reasoning.
{minted} [ frame=lines, numbersep=2mm, breaklines=true, fontsize=, highlightlines=3, highlightcolor=yellow, linenos, tabsize=2, breaklines, ] diff void foo(int input) int n = input * 23; if (3465>=n>=2287) //dangerous code need to be tested
@@
{minted} [ frame=lines, numbersep=2mm, breaklines=true, fontsize=, highlightlines=3,5,7,8, highlightcolor=yellow, linenos, tabsize=2, breaklines, ] diff void bar (int nbits) FFTContext *s = malloc(sizeof (* s)); if (s & nbits==100)) free( s); else … return s;
@@
To construct CodeSense , we collected 544 Python, 100 C, and 100 Java projects. To enable supervised evaluation, we develop a framework that automatically extracts ground-truth annotations for these tasks using static and dynamic program analysis. We build, execute and log run-time execution values and traces of the real-world projects. Our dataset includes a total of 2125 Python, 876 C, and 875 Java unique functions, based on which, we curated 4495 samples with their ground truth in our benchmark. Please see more details in Section 2.
Using our benchmark, we evaluated 14 state-of-the-art (SOTA) LLMs and investigated six research questions regarding the models’ code semantics reasoning capabilities. Previous work has shown that models did not perform well on code reasoning tasks such as input/output prediction [8] and vulnerability detection [11]; to understand why models fail and identify places for improvement, we investigate: RQ1: Does increasing code size make semantic reasoning more difficult? RQ2: Which types of program statements are easier or harder for models to reason about? RQ3: How do models perform on code properties critical for SE tasks, such as predicting pointer aliasing, loop iteration counts, and branch conditions? RQ4: How effective are different prompting strategies in improving semantic reasoning? RQ5: Can models reason approximately when exact values or semantics are hard to infer? RQ6: Do models perform better on some programming languages than others?
Our results reveal that current LLMs, including SOTA models like Claude 3.5  [12], GPT-4o-mini  [13], and Gemini 1.5  [14] struggle with fine-grained code semantics. They often fail to reason about even single statements from real-world code—particularly arithmetic expressions and API calls—and perform poorly on tasks involving loop values and iteration counts. Basic chain-of-thought prompting offers limited benefit, and few-shot prompting yields only modest improvements. In-context learning is most effective when prompts define new concepts or include highly relevant examples. Interestingly, models can correlate natural language code semantics questions with certain code patterns. For instance, when the code contains assignments like p = q, models correctly respond to the prompt "do p and q at <line> alias the same memory address" even in zero-shot settings. Similarly, models reliably infer loop bounds in explicit cases such as for i in range(100):. Among the 14 models evaluated, Claude 3.5 consistently achieved the best performance. We also observe that Java and Python code are generally easier for models to reason about than C, and that input prediction (i.e., reverse semantic inference) remains among the most challenging tasks.
@@ 2.1 Defining a Spectrum of Code Reasoning Tasks
We design fine-grained code reasoning tasks, aiming to (1) help pinpoint models’ code reasoning challenges at a fine-granularity, (2) help models to improve a variety of SE applications that are linked to these fine-grained reasoning tasks.
@@
Task 4: Approximation of code semantics (RQ5): Reasoning about concrete values for the above tasks is very challenging. Sometimes, to solve an SE task, we may only need an approximate value of code semantics. For example, in Fig. 1(a), the models do not have to generate a concrete number like input=120; it is sufficient for models to tell us that an integer input between 100-150 can trigger the dangerous code. We designed a set of abstract values for different data types, following prior literature [9] and evaluate if the models can predict abstract values correctly.
@@ 2.2 Collecting and Tracing Real-world Multi-Lingual Software Projects
Constructing ground truth for the set of code-semantics tasks is a great challenge. We collected a total of 744 real-world software projects of Python, C and Java from GitHub. We developed a framework and tool chain to build the projects, run tests and collect the execution traces which contain values, data types, names of function calls and memory addresses (for C code) at each statement. We developed analysis tools to extract ground truth for the benchmark tasks from those fine-grained code semantics data.
Python. We collected 1489 GitHub repositories from the PyPIbugs dataset  [15]. We removed projects that don’t contain test cases or have not been updated in the last four years, and obtained 544 projects. We first installed dependencies for each project and used pytest  [16] to run tests, and Pysnooper  [17] for tracing.
C. We used 100 projects curated in OSS-Fuzz [18]. We built and fuzzed the real-world projects using the OSS-Fuzz infrastructure in the docker environment with project-wise fuzzing harnesses. We developed a tracing framework built on the GNU debugger (GDB)  [19].
Java. We collected 100 projects from the SF110 dataset [20]. We used EvoSuite [21] to generate and run test cases and developed our tracing tool on top of Java Debugger [22] to record the execution details of the projects.
@@ 2.3 Data Filtering
We collected whole program traces, from which we curated unique functions based on their entry and exit points in the execution trace logs. We excluded functions that only contains comments, too lengthy to fit into the models’ context, and the functions which don’t have meaningful functionality in their body. We obtained a total of 2125 Python, 876 C and 875 Java unique functions, with the sizes ranging from 3 to 516 lines of code. From these unique functions, we curated our task-specific datasets.
In real-world code, we face many complexities, e.g., the input of a function and the values of a variable can be complex types. As the first step of probing models to reason about fine-grained code semantics for real-world code, we focus on ground truth values of primitive data types in all tasks, including int, float, str, bool, list, pointer, double, dictionary, tuple, etc. In the evaluation, we show that even for values of primitive types, the models face challenging to predict them. We collected a total of 4495 samples from Python, C and Java, and constructed the ground truth for the above tasks and used for evaluation. See Table 2.
@@
Task 2/Task 4: Statement 545 497 1042
@@ 3 Evaluation
We designed five different natural language prompt templates (see Appendix/data package), and ran them on a sampled dataset for each model. We observed that prompt templates are model-sensitive, but not task-sensitive. So we select a template for each model for all the tasks. We prompted the models to give a response inside specific tags (<ans> </ans>) and considered the response inside that tag to compare with the ground truth, as done in  [8]. For our evaluation metrics, we used accuracy (exact matching of the generated outputs of the models and the ground truth label).
@@ 3.2 Results for RQ2: Statement-level code semantics
We observe that arithmetic and API/calls are the most statement types, even for the best reasoning models like GPT4.0-mini and Claude 3. For APIs, we sampled frequently used third-party libraries, like os, sys, time and math installed by pip, but the models do not have knowledge about their execution semantics. Models handle better for predicting Boolean values such as the output of a comparison statement, and also statements where constant is assigned to a variable. The models may understand the assignment operator "=" and have captured easy patterns like "a=3 indicates a has value 3 after executing this statement". We did not observe significant advantage of reasoning models over non-reasoning models for this task.
@@ 3.5 Results for RQ5: Approximation of code semantics
In Fig. 7, we observed that the models reported better performance to predict an approximation of code semantics, for both statement (Fig. 7(b)) and loop (Fig. 7(c)) predictions. Interestingly, when we provide only the definition of "abstract" values (mapping from a range of concrete values to an abstract value) in the prompt, without giving an example showing an "abstract" output for a given input, the models cannot predict abstract values better than concrete values (Fig. 7(a)). Most models failed to apply definitions directly to the query examples; however, when we provide 3-shots of examples in the prompt, all the models can predict abstract values better than concrete ones.
@@ 4 Related Work
Code Reasoning Benchmarks. CruxEval [8] assesses LLMs’ performance on synthetic Python programs using the task of input-output prediction for a function. CruxEval-X [23] extends this work to multilingual settings by translating synthetic Python programs in CruxEval to other languages using LLMs. REval [24] evaluated branch prediction tasks using ClassEval and HumanEval [25]. CodeMind [26] proposed output prediction and code synthesis tasks on existing code benchmarks [27, 28, 8, 29]. They found that LLM reasoning capabilities deteriorate as program complexity increases [26, 30]. Our benchmark CodeSense is the first that used real-world Python, C and Java code to evaluate LLMs’ reasoning capabilities. While most code reasoning benchmarks reason about function level execution semantics, we proposed and made ground truth available for a spectrum of fine-grained reasoning tasks regarding program behaviors within a function.
Other Code Application Benchmarks. SWE-Bench [5] used the task of generating patches to resolve a given GitHub issues for real-world Python projects. SWE-PolyBench [6] extends this work to other programming languages. KGym [7] delivered a benchmark consisting of Linux kernel crash data and evaluated LLMs’ capabilities of resolving Linux kernel crashes. These benchmarks focus on task-specific performance rather than fine-grained code semantics understanding. There are also benchmarks for code generation [31, 32, 33, 34, 35] and code completion [36, 37]. However, most of these datasets—such as BigCodeBench [3] and CodeBenchGen [4] are restricted to a single language (primarily Python) and extracted from isolated competitive programming problems.
@@ A.1 Computation Resources and Inference Tools
  • Software Stack: vLLM  [38] for optimised transformer inference
@@ A.3.1 Real World Project Collection
    Python Project Collection: To collect real-world Python projects, we adopted two approaches. 1) We cloned all 1489 repositories from GitHub that appear in the PyPIBugs dataset, which was released in 2021 [15]. 2) To avoid missing popular projects after 2021, we use GitHub API to search for repositories that are marked as mainly written in Python and get the results according to the descending order of the number of stars. To maximise the probability that we can execute them easily with the pytest module, we only consider projects that seemingly have a testing folder at the top or second level. For better compatibility and the reflection of the recent trend of programming styles, we further filtered out projects that have not been updated in the last four years. Finally, we got 544 projects.
@@
    Java Dataset Collection: For Java, we aimed to gather a diverse set of real-world projects to have a comprehensive trace analysis, which aligns with our trace dataset generation objective. We have used EvoSuite [21] for test suite generation, and the SF110 dataset has been used as it is recommended by EvoSuite. This choice ensures compatibility and a high testing coverage rate.
@@ C OSS-Fuzz (100) Fuzzing GDB
A.4 Detailed Description of Tasks

### v2

@@ 1 Introduction
Recent years have witnessed the emergence of numerous benchmarks for evaluating coding-related tasks. However, the majority of these efforts have focused on code generation using synthetic or narrowly scoped data—for example, HumanEval+ (Liu et al., 2023), LiveCodeBenchmark (Jain et al., 2024), Bigcodebench (Zhuo et al., 2024), and CodeBenchGen (Xie et al., 2024)—often extracted from isolated competitive programming problems. Consequently, they fail to capture the complexity and structure of real-world software development. Other benchmarks that incorporate real-world code, such as SWE-Bench (Jimenez et al., 2024), SWE-PolyBench (Rashid et al., 2025), and KGym (Mathai et al., 2024), tend to evaluate only task-specific performance (e.g., patch generation for GitHub issues), making it difficult to assess whether models exhibit generalizable semantic understanding. Finally, reasoning-focused benchmarks, such as CruxEval (Gu et al., 2024), primarily target function-level input/output prediction over short and synthetic code fragments involving random string operations. Such settings neglect the fine-grained semantic reasoning about internal program behavior and properties, data dependencies, and control structures required to solve a variety of SE tasks for complex real-world software systems.
To this end, we propose CodeSense, a benchmark for fine-grained code semantic reasoning, constructed from real-world GitHub projects in Python, C, and Java (see Table 1). CodeSense  introduces a spectrum of reasoning tasks at statement, code-block, and function levels, targeting essential semantic properties frequently needed across SE activities. For example, predicting loop iteration counts is critical for input/output prediction, performance analysis, and detecting infinite loops (e.g., denial-of-service vulnerabilities). Branch condition prediction and reasoning about pointers in C code are important for test input generation and memory safety assurance. As illustrated in Figures 1(a) and 1(b), fine-grained semantic reasoning about arithmetic operations, control flow, and API semantics is foundational for a variety of SE applications. Prior work (Ding et al., 2023a; 2024) has shown that incorporating semantic signals during training improves model performance on code generation, branch prediction, code clone detection, program repair and vulnerability detection tasks, motivating our design of CodeSense to comprehensively evaluate models’ capabilities for semantic reasoning. Their experimental results showed that even using dynamic values collected from small implementations (<100 lines of code), the models are able to improve downstream tasks for real-world code. Similarly, Valgrind  (Nethercote and Seward, 2007), a widely used tool and deployed for more than 20 years (Nethercote, 2022), uses dynamic values for reasoning tasks, such as detecting and debugging memory errors, and optimizing programs.
void foo(int input){ int n = input * 23; if (3465>=n>=2287){ //dangerous code need to be tested } }
@@
void bar (int nbits) { FFTContext *s = malloc(sizeof (* s)); if (s && nbits==100)) free(& s); else ... return s; }
@@
Using our benchmark, we evaluated 14 state-of-the-art (SOTA) LLMs and investigated six research questions regarding the models’ code semantics reasoning capabilities. Previous work has shown that models did not perform well on code reasoning tasks such as input/output prediction (Gu et al., 2024) and vulnerability detection (Steenhoek et al., 2025); to understand why models fail and identify places for improvement, we investigate: RQ1: Does increasing code size make semantic reasoning more difficult? RQ2: Which types of program statements are easier or harder for models to reason about? RQ3: How do models perform on code properties critical for SE tasks, such as predicting pointer aliasing, loop iteration counts, and branch conditions? RQ4: How effective are different prompting strategies in improving semantic reasoning? RQ5: Can models reason approximately when exact values or semantics are hard to infer? RQ6: Do models perform better on some programming languages than others?
Our results reveal that current LLMs, including SOTA models like Claude 3.5 (Anthropic, 2024), GPT-4o-mini  (OpenAI, 2024), and Gemini 1.5  (Google, 2025) struggle with fine-grained code semantics. They often fail to reason about even single statements from real-world code—particularly arithmetic expressions and API calls—and perform poorly on tasks involving loop values and iteration counts. Basic chain-of-thought prompting offers limited benefit, and few-shot prompting yields only modest improvements. In-context learning is most effective when prompts define new concepts or include highly relevant examples. Interestingly, models can correlate natural language code semantics questions with certain code patterns. For instance, when the code contains assignments like p = q, models correctly respond to the prompt "do p and q at <line> alias the same memory address" even in zero-shot settings. Similarly, models reliably infer loop bounds in explicit cases such as for i in range(100):. Among the 14 models evaluated, Claude 3.5 consistently achieved the best performance. We also observe that Java and Python code are generally easier for models to reason about than C, and that input prediction (i.e., reverse semantic inference) remains among the most challenging tasks.
@@ 2.1 Defining a Spectrum of Code Reasoning Tasks
To design tasks for evaluating LLMs’ capabilities of code semantic reasoning, we first considered the definition of code semantics. In programming languages and software engineering, code semantics —“what is the meaning of this code” — are defined as what is the output value given the input of a code snippet. Such fine-grained reasoning tasks are directly related to end-tasks in software engineering. For example, previous work (Ding et al., 2023a) shows that when fine-tuned with statement-level values, the performance of the models improved for vulnerability detection, branch prediction and code clone detection. Prior study (Steenhoek et al., 2025) reported that although recent LLMs improved math reasoning and natural language reasoning significantly, they are still insufficient for handling end-tasks related to code reasoning. To help locate the weakness of models’ code reasoning at a fine-granularity and help models to improve a variety of SE applications that are linked to the fine-grained reasoning steps, we designed the following code reasoning tasks:
@@
Task 4: Approximation of code semantics (RQ5): Reasoning about concrete values for the above tasks is very challenging. Sometimes, to solve an SE task, we may only need an approximate value of code semantics. For example, in Fig. 1(a), the models do not have to generate a concrete number like input=120; it is sufficient for models to tell us that an integer input between 100-150 can trigger the dangerous code. We designed a set of abstract values for different data types, following prior literature (Ding et al., 2023a) and evaluate if the models can predict abstract values correctly.
@@ 2.2 Collecting and Tracing Real-world Multi-Lingual Software Projects
Constructing ground truth for the set of code-semantics tasks is a great challenge. We collected a total of 744 real-world software projects of Python, C and Java from GitHub. We developed a framework and tool chain to build the projects, run tests and collect the execution traces which contain values, data types, names of function calls and memory addresses (for C code) at each statement. We developed analysis tools to extract ground truth for the benchmark tasks from those fine-grained code semantics data. Please check our Appendix  A.4 for our language selection rationale and how our framework can be easily extended to other languages.
Python. We collected 1489 GitHub repositories from the PyPIbugs dataset (Allamanis et al., 2021). We removed projects that don’t contain test cases or have not been updated in the last four years, and obtained 544 projects. We first installed dependencies for each project and used pytest (Krekel et al., 2004) to run tests, and Pysnooper  (Rachum et al., 2019) for tracing.
C. We used 100 projects curated in OSS-Fuzz (Arya et al., 2023). We built and fuzzed the real-world projects using the OSS-Fuzz infrastructure in the docker environment with project-wise fuzzing harnesses. We developed a tracing framework built on the GNU debugger (GDB)  (Free Software Foundation, ).
Java. We collected 100 projects from the SF110 dataset (Fraser and Arcuri, 2012). We used EvoSuite (Fraser and Arcuri, 2011) to generate and run test cases and developed our tracing tool on top of Java Debugger (Oracle Corporation, 2025) to record the execution details of the projects.
It should be noted that although our repository are in the training dataset. However, the models are asked to predict dynamic information, which we generated by running test inputs over programs, using tools like fuzzers. Those values are not in the repos and have never been seen by the models. In fact, this is the advantage of our benchmark. The future benchmark designers can use our testing and tracing tools to run other inputs to freshly generate more data that are new to the models.
@@ 2.3 Data Filtering
We collected whole program traces, from which we curated unique functions based on their entry and exit points in the execution trace logs. We excluded functions that only contains comments, too lengthy to fit into the models’ context, and the functions which don’t have meaningful functionality in their body. For example, some functions only contain one statement like “return 0”, or “printf(“…”) or some functions are just a wrapper for another function which we have tested, such as "void myfunc(){ func();}". We obtained a total of 2125 Python, 876 C and 875 Java unique functions, with the sizes ranging from 3 to 516 lines of code. From these unique functions, we curated our task-specific datasets.
In real-world code, we face many complexities, e.g., the input of a function and the values of a variable can be complex types. As the first step of probing models to reason about fine-grained code semantics for real-world code, we focus on ground truth values of primitive data types in all tasks, including int, float, str, bool, list, pointer, double, dictionary, tuple, etc. In the evaluation, we show that even for values of primitive types, the models face challenging to predict them. We collected a total of 4483 samples from Python, C and Java, and constructed the ground truth for the above tasks and used for evaluation. See Table 2.
@@
Task 2/Task 4: Statement 545 485 1030
@@ 3 Evaluation
We designed five different natural language prompt templates (see Appendix/data package), and ran them on a sampled dataset for each model. We observed that prompt templates are model-sensitive, but not task-sensitive. So we select a template for each model for all the tasks. We prompted the models to give a response inside specific tags (<ans> </ans>) and considered the response inside that tag to compare with the ground truth, as done in  Gu et al. (2024). For our evaluation metrics, we used accuracy (exact matching of the generated outputs of the models and the ground truth label).
@@ 3.2 Results for RQ2: Statement-level code semantics
We observe that arithmetic and API/calls are the most statement types, even for the best reasoning models like GPT4.0-mini and Claude 3. For APIs, we sampled frequently used third-party libraries, like os, sys, time and math installed by pip, but the models do not have knowledge about their execution semantics. We also experimented with adding the API definitions in the prompt, but it didn’t increase the performance significantly A.9. Models handle better for predicting Boolean values such as the output of a comparison statement, and also statements where constant is assigned to a variable. The models may understand the assignment operator "=" and have captured easy patterns like "a=3 indicates a has value 3 after executing this statement". We did not observe significant advantage of reasoning models over non-reasoning models for this task.
@@ 3.5 Results for RQ5: Approximation of code semantics
In Fig. 7, we observed that the models reported better performance to predict an approximation of code semantics, for both statement (Fig. 7(b)) and loop (Fig. 7(c)) predictions. See also Table 7 in Appendix A.8.3 for comparing with random baselines. Interestingly, when we provide only the definition of "abstract" values (mapping from a range of concrete values to an abstract value) in the prompt, without giving an example showing an "abstract" output for a given input, the models cannot predict abstract values better than concrete values (Fig. 7(a)). Most models failed to apply definitions directly to the query examples; however, when we provide 3-shots of examples in the prompt, all the models can predict abstract values better than concrete ones.
@@ 4 Related Work
Code Reasoning Benchmarks. CruxEval Gu et al. (2024) assesses LLMs’ performance on synthetic Python programs using the task of input-output prediction for a function. CruxEval-X Xu et al. (2024) extends this work to multilingual settings by translating synthetic Python programs in CruxEval to other languages using LLMs. REval Chen et al. (2024) evaluated branch prediction tasks using ClassEval and HumanEval Chen et al. (2021a). CodeMind Liu et al. (2024) proposed output prediction and code synthesis tasks on existing code benchmarks Chen et al. (2021b); Austin et al. (2021); Gu et al. (2024); Puri et al. (2021). They found that LLM reasoning capabilities deteriorate as program complexity increases Liu et al. (2024); Zhang et al. (2024a). Our benchmark CodeSense is the first that used real-world Python, C and Java code to evaluate LLMs’ reasoning capabilities. While most code reasoning benchmarks reason about function level execution semantics, we proposed and made ground truth available for a spectrum of fine-grained reasoning tasks regarding program behaviors within a function.
Other Code Application Benchmarks. SWE-Bench Jimenez et al. (2024) used the task of generating patches to resolve a given GitHub issues for real-world Python projects. SWE-PolyBench Rashid et al. (2025) extends this work to other programming languages. KGym Mathai et al. (2024) delivered a benchmark consisting of Linux kernel crash data and evaluated LLMs’ capabilities of resolving Linux kernel crashes. These benchmarks focus on task-specific performance rather than fine-grained code semantics understanding. There are also benchmarks for code generation Li et al. (2024); Zhang et al. (2024b); Yu et al. (2024); Chen et al. (2025); Du et al. (2023) and code completion Izadi et al. (2024); Ding et al. (2023b). However, most of these datasets—such as BigCodeBench Zhuo et al. (2024) and CodeBenchGen Xie et al. (2024) are restricted to a single language (primarily Python) and extracted from isolated competitive programming problems.
@@ A.1 Computation Resources and Inference Tools
  • Software Stack: vLLM  Kwon et al. [2023] for optimised transformer inference
@@ A.3.1 Real World Project Collection
    Python Project Collection: To collect real-world Python projects, we adopted two approaches. 1) We cloned all 1489 repositories from GitHub that appear in the PyPIBugs dataset, which was released in 2021 Allamanis et al. [2021]. 2) To avoid missing popular projects after 2021, we use GitHub API to search for repositories that are marked as mainly written in Python and get the results according to the descending order of the number of stars. To maximise the probability that we can execute them easily with the pytest module, we only consider projects that seemingly have a testing folder at the top or second level. For better compatibility and the reflection of the recent trend of programming styles, we further filtered out projects that have not been updated in the last four years. Finally, we got 544 projects.
@@
    Java Dataset Collection: For Java, we aimed to gather a diverse set of real-world projects to have a comprehensive trace analysis, which aligns with our trace dataset generation objective. We have used EvoSuite Fraser and Arcuri [2011] for test suite generation, and the SF110 dataset has been used as it is recommended by EvoSuite. This choice ensures compatibility and a high testing coverage rate.
@@ C OSS-Fuzz (100) Fuzzing GDB
A.4 Language Selection Rationale and Extensibility
The three programming languages in our benchmark are important and representative for programming language features and real-world applications: C is a low level programming language and useful for building systems, Java has object-oriented programming features, are widely used for building enterprise and web applications, Python is important for data science and AI applications. Our benchmark is extensible, third-parties (as well as ourselves in future) can add more languages. Our scripts, prompts and methodologies can be adapted for new programming languages to (1) select and download programs of a programming language from GitHub, (2) fuzz for generating test inputs, (3) trace and curate ground truth data (4) provide input and parse output when interacting with models. Instead of GDB (for C), Pysnooper (for Python) and Java Debugger Oracle Corporation (for Java), we will need to plug in debugging tools for new programming languages.
You may ask “we have compilers and code execution tools, why do we need models to predict dynamic information”—- Predicting dynamic values is not only useful for executing programs, but required for many other downstream tasks. For example, in Fig. 1, to generate test input that can exercise a true branch, the models need to know how each operator in statements updates the values. The key difference is that: when using code execution tools, we give one input and ask for the output, but in other downstream tasks, we require models to first understand fine-grained semantics and then find inputs that can satisfy certain constraints. Even for compiling and executing programs, we may benefit from LLMs prediction, as compilation and execution can be time-consuming and hard to be configured, especially for legacy code. This ability of LLMs is particularly important when the user cannot run the code snippet, e.g., missing dependencies or unavailable resources. Predicting input/output values as code reasoning tasks have been established by prior research [Gu et al., 2024], [Chen et al., 2021b], [Yan et al., 2024]. Our work extended prior research by introducing real-world projects and fine-grained tasks supported only by our tracing framework.
A.5 Detailed Description of Tasks